OpenStudy (anonymous):

Determine if the following series is convergent or divergent. sum of ( (sin n) / n)^2 from n = 1 to n = inf not sure it should be here though... I can't find Calculus 2 room. Thanks.

OpenStudy (anonymous):

Easier look of this problem : \[\sum_{n = 1}^{\infty} \ (\frac{ \sin n }{ n }) ^ {2}\]

OpenStudy (kainui):

I thought I was in a wormhole to the past, but then I realized I was in a wormhole to the past.

OpenStudy (anonymous):

lol .... I also think the answer is converge (to zero) as same as yesterday one ..... but I don't know how to prove it.

OpenStudy (anonymous):

DAMNIT .... Got it.... because \[\sum_{n = 1}^{\infty} (\frac{ \sin n }{ n }) ^{2} \le \sum_{n = 1}^{\infty} (\frac{ 1 }{ n }) ^{2}\] for every n because \[0 \le \sin n \le 1\] for every n and the second series converge because it is p-serie. so the first series converge. Thank you sir !!!!